# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v4

@@ 3.4. New black hole solutions
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}+{4\over(\rho_{+}^{2}-\rho_{-}^{2})}\left[d{\widetilde{x}}+{(\rho_{+}^{2}-\rho_{-}^{2})(1-2\lambda)\over 4l}d{\widehat{t}}\right]^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|.}$ (3.48)3.48( 3.48 )
We note that for $\rho_{-}=0$ it is equivalent to the solution (3.27) in the sense both have a horizon at $\rho_{+}$ and a singularity at $\rho=\rho_{-}=0$. However, here the black hole is in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field.
The black hole solutions given in (3.44) are described by the four parameters $\rho_{-}$, $\rho_{+}$, $\gamma$ and $\lambda$. Thus, in general they are not identical. To the best of my knowledge the black hole solutions (3.44) have not been reported elsewhere in the literature. We next find the (analogous) ADM masses and angular momenta of the solutions (3.25) and (3.44). The ADM masses are given in a similar way as in the BTZ black hole by the asymptotic values of the Brown-York quasi-local masses at radial infinity. The angular momenta are given by the Brown-York quasi-local angular momenta which (as we momentarily show) are independent of $\rho$ and thus constant.
@@ 4. Dilatonic rotating black hole solutions
In this section we consider the black hole solutions (3.25) and (3.44). They are exact solutions of classical string theory, i.e., they have exact worldsheet CFT descriptions. The reason is essentially because we found them by applying T-dualities and coordinate transformations on $AdS_{3}$. More on the deformed CFTs in a future work. The solutions are again given below for convenience. The rotating black hole solution (3.25) is given by
@@
$\eqalign{\widetilde{R}&=e^{4\phi}\left[R-4\nabla^{2}\alpha-2(\nabla\alpha)^{2}\right],\cr&=-{(\rho_{+}^{2}-\rho_{-}^{2})^{2}\over 4l^{2}}{\left[\rho_{+}^{2}-\rho_{-}^{2}+2(\rho^{2}-\rho_{-}^{2})\right]\over(\rho^{2}-\rho_{-}^{2})^{3}}.}$ (4.9)4.9( 4.9 )
@@
$\eqalign{d\widetilde{s}^{2}&=N^{2}(r)dr^{2}-K^{2}(r)dt^{2}+r^{2}\left(d\theta+Zdt\right)^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}f(r)|,}$ (4.14)4.14( 4.14 )
@@
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=2\rho_{+}$. The metric components and dilaton $N,K,Z$ and $f$ are some known functions of $r$.
@@
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=(1-\gamma)\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}$. The metric components and dilaton $N,K,Z$ and $f$ are some other known functions of $r$.
The black hole solution (4.13), as we mentioned earlier, is already known in the literature and was first obtained earlier in [7]. The authors [7] studied the Einstein-scalar gravity in three spacetime dimensions with the scalar in an exponential potential. The solution (4.13) belongs to the family of black hole solutions obtained in [7]. In [7] it corresponds to the case $N=1$. To write (4.13) in the form given in [7] we need to make the change of coordinate
@@
where $\varepsilon$ determines the zero of the energy. It is given by $\varepsilon=LK^{\prime}$ for a particular solution. Primes denote derivatives with respect to $r$. $j$ agrees with the usual Komar angular momentum formula (adapted for $(2+1)$ dimensions). For the derivation of (4.20) and review see the appendix. See also, e.g., [[7]7,,  [31]31,,  [32]32,,  [33]33].
@@
It is important to keep in mind that the mass $m$ is measured by the static observer whose four-velocity is $u^{\mu}_{(1)}=\delta^{\mu}_{t}=(1,0,0)$, i.e., tangent to the static Killing field $\xi^{\mu}=(\partial_{t})^{\mu}$. The energy $E$ is measured by the stationary observer whose four-velocity is $u^{\mu}_{(2)}=(1/N)\delta^{\mu}_{t}-(Z/N)\delta^{\mu}_{\phi}$.19 Note in general $u^{\mu}_{(2)}$ gives, in the large $r$ limit, the Kodama field [[34]34,,[35]35]. The Kodama vector field is given by $\eqalign{k&=k^{t}\partial_{t}+k^{\phi}\partial_{\phi},\cr k^{t}&={1\over N}LK^{\prime},\cr k^{\phi}&=-{Z\over N}LK^{\prime}+{1\over 2K^{2}}\cdot{LK^{3}Z^{\prime}\over N},\cr&=-{Z\over N}LK^{\prime}+{j\over 2K^{2}}.}$ (4.22)4.22( 4.22 ) In general, the Kodama field is not a Killing field. At large $r$ it is proportional to $Nu^{\mu}_{(2)}=(\partial_{t})^{\mu}-Z(\partial_{\phi})^{\mu}$. Its world-line is perpendicular to the $t={\tenrm constant}$ space-like hyper-surface. Thus, the energy and mass are equal only when the shift $Z=0$ and the lapse $N=1$.
@@
This solution (4.26) gives $\varepsilon=\sqrt{-\Lambda\rho^{2}}$. The particular solution is chosen such that at constant time or $\varphi$ it asymptotically coincides or merges at large $\rho$ smoothly with the solution (2.6) [[3]3,,[4]4]. Thus, the quasi-local mass $m$ is given by20 Note here the boundary $B_{t}$ is at large $\rho$, see Appendix C. Thus, in the expressions for $m$ and $E$, $\rho$ is assumed to be large. The same applies for the quasi-local masses and energies that appear below. Here they serve as intermediate results.
@@
Note the quais-local mass $m$ depends on $\varepsilon$, i.e., the reference background. Different $\varepsilon$ gives different $m$. The global $AdS_{3}$ (2.3), for instance, gives $\varepsilon=\sqrt{-\Lambda\rho^{2}+1}$. However, the ADM mass $M$ is independent of the reference solution or regularization and it is finite.
@@
Similarly, for the black hole solutions (4.16) we find that the angular momenta are given by21 Here we set $g_{s}=1$.
$J=2\left({1-\gamma^{2}\over l}\right)=Q,\quad-1\leq\gamma\leq 0.$ (4.38)4.38( 4.38 )
@@
At $\lambda=-1/2\gamma$ we note that the Kalb-Ramond field is zero at infinity. The ADM mass and angular momentum are
@@
A solution with a non-zero asymptotic Kalb-Ramond field is thus equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity. In particular, there is dragging, usually known as the Lense-Thirring effect, at infinity, i.e., $Z(r=\infty)\neq 0$. Thus, locally zero angular momentum observers (ZAMOs) [[36]36,,  [37]37] located at asymptotic infinity rotate with a non-zero constant angular velocity $\Omega=Z(r=\infty)$.
Although non-zero $\Omega$ can certainly lower (and raise) the mass, see e.g. the general expression (C.27) of the mass in the appendix, it does not always result in negative mass solutions. A simple example is the BTZ solution with non-vanishing rotation at asymptotic infinity. To allow rotation at infinity we shift $Z$ in (4.29) by a non-zero constant. We denote the constant by $\overline{\lambda}$ and make the change
@@ 5. Discussion
The black string solutions are described by (3.15) which we again write below for convenience. They are obtained from ${\cal A}_{3}$ by applying the simple coordinate transformation (3.13) (and making the change of variable (3.14)). They carry the axion charge per unit length $Q$ (3.19). The black string solutions are
@@
$\eqalign{e^{-2\phi}&={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}},}$ (5.4)5.4( 5.4 )
$\lambda$ is an arbitrary constant which determines the value of the $B$ field, i.e. $B_{tx}$, at $\rho=\infty$. It does not enter in $H$. However, it has a non-trivial implication. The black string solution (5.3) has inner and outer horizons. This is also a general feature of rotating black holes. The inner horizon is at $\rho=\rho_{-}$ and the outer horizon is at $\rho=\rho_{+}$. In this paper, we in general assume, unless stated otherwise, $x$ is compact and $\rho$ is positive. In the case there is no curvature singularity we identify $x$ so that there is no conical singularity. Its period does not depend on the gauge parameter $\lambda$. To the best of my knowledge the black string solutions (5.3) have not been reported elsewhere in the literature.
For positive $\gamma$, i.e., $0<\gamma\leq 1$, the solution has a curvature singularity at $\rho=\rho_{0}\geq\rho_{+}$ (4.1). Therefore, it cannot be continuously contracted or deformed to the origin without crossing the horizon(s). For $\gamma=1$ the outer horizon becomes singular, see (3.33). The singularity is behind the horizon in the interior region, i.e., $\rho_{0}>\rho_{-}$. Note the interior region is defined in the usual way as the part of the spacetime that contains the singularity. See the Penrose diagrams in Fig. 4 and 5.
@@
The Kalb-Ramond field is invariant under the Lorentz boost. The resulting metric however now has no cross term and thus static, see (3.37). We subsequently applied T-duality transformation along the $\widehat{x}$ direction. This resulted in the black hole solution (3.44),
@@
To the best of my knowledge this black hole solution has not been reported elsewhere in the literature. The black hole solution has a singularity at $\rho=\rho_{-}$ and an event horizon at $\rho=\rho_{+}$. See the Penrose diagram in Fig. 3. Its (analogous) ADM mass and angular momentum are given by (4.38), (4.41)
@@
At $\gamma=-1$ we note that $J=0$ and the mass $M$ is independent of the gauge parameter $\lambda$. $M$ is determined by the radii $\rho_{+}$ and $\rho_{-}$. At $\lambda=-1/2\gamma$, the angular velocity at infinity is zero, i.e., $\Omega=0$. In this case the ADM mass and angular momentum reduce to (4.43)
@@
In this paper we also noted that a solution with a non-zero asymptotic Kalb-Ramond field is equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity. In general, non-zero rotation allowes negative mass solutions.
In an upcoming work we study the thermodynamics and partition functions of the black hole and black string solutions. We also study the casual structure of the solutions in detail.
In this paper we mainly presented black hole and black string solutions. In a future work we hope to study in greater detail the marginal operator that generates the deformation on the worldsheet.
@@
The scalar curvature of the solution (which is described by (5.13) and (5.14)) is $R=-4\xi^{2}\cdot(\rho_{+}^{2}/l^{2})\cdot(1/|\rho^{2}-\rho_{-}^{2}|)$, where $\xi^{2}=1-(\rho_{-}/\rho_{+})^{2}$. The Kretschmann scalar is $K=R^{2}$. Thus, the solution has the peculiar feature that it contains a ring curvature singularity at $\rho={\overline{\rho}}_{+}=\rho_{-}$ in between the inner and outer horizons. The inner horizon is at $\rho={\overline{\rho}}_{-}$. The outer horizon is at $\rho=\rho_{+}$. Note they are not related by sign reflection, i.e. ${\overline{\rho}}_{-}\neq-\rho_{+}$. The Kalb-Ramond field is a constant (field). It can be fixed to zero. The Penrose diagram of the solution is given in Fig. 6. It is obtained by combining Fig. 3 and 5.
Fig. 6: The plot depicts the global structure of the string solution described by (5.13) and (5.14). The solution has a singularity at $\rho=\rho_{-}={\overline{\rho}}_{+}$. It is represented by the zigzag lines.
Note usually the diagram ends at $\rho={\overline{\rho}}_{+}=\rho_{-}$, however here, since there are no closed time-like curves, the diagram extends outward to infinity and inward to zero. Therefore, the maximal extension simply extends infinitely in both the vertical directions.
We expect the black hole solution (described by (5.13) and (5.14)) to have a generalization in (3 + 1) dimensions with similar feature. That is, with a curvature singularity in between the inner and outer horizons. We hope to study the corrections in $g_{s}$ and their roles in resolving the singularity in a future work.

### v5

@@ 3.4. New black hole solutions
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}+{4\over(\rho_{+}^{2}-\rho_{-}^{2})}\left[d{\widetilde{x}}+{(\rho_{+}^{2}-\rho_{-}^{2})(1-2\lambda)\over 4l}d{\widehat{t}}\right]^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|.}$ (3.56)3.56( 3.56 )
We note that for $\rho_{-}=0$ it is equivalent to the solution (3.34) in the sense both have a horizon at $\rho_{+}$ and a singularity at $\rho=\rho_{-}=0$. However, here the black hole is in a spacetime with non-zero angular velocity at radial infinity and no $B$ field.
The black hole solutions given in (3.51) are described by the four parameters $\rho_{-}$, $\rho_{+}$, $\gamma$ and $\lambda$. Thus, in general they are not identical. To the best of my knowledge the black hole solutions (3.51) have not been reported elsewhere in the literature. We next find the (analogous) ADM masses and angular momenta of the solutions (3.32) and (3.51). The ADM masses are given in a similar way as in the BTZ black hole by the asymptotic values of the Brown-York quasi-local masses at radial infinity. The angular momenta are given by the Brown-York quasi-local angular momenta which (as we momentarily show) are independent of $\rho$ and thus are constant.
@@ 4. Dilatonic rotating black hole solutions
In this section we consider the black hole solutions (3.32) and (3.51). They are exact solutions of classical string theory, i.e., they have exact world-sheet CFT descriptions. The reason is essentially because we found them by applying T-dualities and coordinate transformations on $AdS_{3}$. More on the deformed CFTs in a future work. The solutions are again given below for convenience. The rotating black hole solution (3.32) is given by
@@
$\eqalign{\widetilde{R}&=e^{4\phi}\left[R-4\nabla^{2}\alpha-2(\nabla\alpha)^{2}\right],\cr&=-{4g_{s}^{4}\over l^{2}}{\left[\rho_{+}^{2}-\rho_{-}^{2}+2(\rho^{2}-\rho_{-}^{2})\right]\over(\rho^{2}-\rho_{-}^{2})^{3}}.}$ (4.9)4.9( 4.9 )
@@
$\eqalign{d\widetilde{s}^{2}&=L^{-2}(r)dr^{2}-N^{2}(r)dt^{2}+r^{2}\left(d\theta+Zdt\right)^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}f(r)|,}$ (4.14)4.14( 4.14 )
@@
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=2\rho_{+}$. The metric components and dilaton, i.e., $N,L,Z$ and $f$, are some known functions of $r$.
@@
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=(1-\gamma)\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}$. The metric components and dilaton, i.e. $N,L,Z$ and $f$, are some other known functions of $r$.
The black hole solution (4.13), as we mentioned in the previous section, is already known in the literature and was first obtained earlier in [7]. The authors [7] studied the minimally coupled Einstein-scalar gravity in three spacetime dimensions with the scalar in an exponential potential. The solution (4.13) belongs to the family of black hole solutions obtained in [7]. In [7] it corresponds (using their notation) to the case $N=1$. To write (4.13) in the form given in [7] we need to make the change of coordinate
@@
where $\varepsilon$ determines the zero of the energy. It is given by $\varepsilon=LK^{\prime}$ for a particular solution. Primes denote derivatives with respect to $r$. $j$ agrees with the usual Komar angular momentum formula (adapted for $(2+1)$ dimensions). For the derivation of (4.20) and review see the Appendix. See also, e.g., [[7]7,,  [34]34,,  [35]35,,  [36]36].
@@
It is important to keep in mind that the mass $m$ is measured by the static observer whose four-velocity is $u^{\mu}_{(1)}=\delta^{\mu}_{t}=(1,0,0)$, i.e., tangent to the static Killing field $\xi^{\mu}=(\partial_{t})^{\mu}$. The energy $E$ is measured by the stationary observer whose four-velocity is $u^{\mu}_{(2)}=(1/N)\delta^{\mu}_{t}-(Z/N)\delta^{\mu}_{\phi}$.25 Note in general $u^{\mu}_{(2)}$ gives, in the large $r$ limit, the Kodama field [[37]37,,[38]38]. The Kodama vector field is given by $\eqalign{k&=k^{t}\partial_{t}+k^{\phi}\partial_{\phi},\cr k^{t}&={1\over N}LK^{\prime},\cr k^{\phi}&=-{Z\over N}LK^{\prime}+{1\over 2K^{2}}\cdot{LK^{3}Z^{\prime}\over N},\cr&=-{Z\over N}LK^{\prime}+{j\over 2K^{2}}.}$ (4.22)4.22( 4.22 ) In general, the Kodama field is not a Killing field. At large $r$ it is proportional to $Nu^{\mu}_{(2)}=(\partial_{t})^{\mu}-Z(\partial_{\phi})^{\mu}$. Its world-line is perpendicular to the $t={\tenrm constant}$ space-like hyper-surface. Thus, the energy and mass are equal only when the shift $Z=0$ and the lapse $N=1$.
@@
This solution i.e., (4.26), gives $\varepsilon=\sqrt{-\Lambda\rho^{2}}$. The particular solution is chosen such that at constant time or $\varphi$ it asymptotically coincides or merges at large $\rho$ smoothly with the solution (2.6) [[3]3,,[4]4]. Thus, the quasi-local mass $m$ is given by26 Note here the boundary $B_{t}$ is at large $\rho$, see Appendix C. Thus, in the expressions for $m$ and $E$, $\rho$ is assumed to be large. The same applies for the quasi-local masses and energies that appear below. Here they serve as intermediate results.
@@
Note the quasi-local mass $m$ depends on $\varepsilon$, i.e., the reference background. Different $\varepsilon$ gives different $m$. The global $AdS_{3}$ (2.3), for instance, gives $\varepsilon=\sqrt{-\Lambda\rho^{2}+1}$. However, the ADM mass $M$ is independent of the reference solution or regularization and it is finite.
@@
Similarly, for the black hole solutions (4.16) we find that the angular momenta are given by
$J=2\left({1-\gamma^{2}\over l}\right)={Q\over g_{s}},\quad-1\leq\gamma\leq 0.$ (4.38)4.38( 4.38 )
@@
At $\lambda=-1/2\gamma$ we note that the Kalb-Ramond field is zero at infinity. The ADM mass and angular momentum are27 Here we set $\kappa_{0}^{2}=\pi$ (in some length unit) and $g_{s}=1$, see (3.8).
@@
A solution with a non-zero asymptotic Kalb-Ramond field is thus equivalent to a solution with no Kalb-Ramond field and non-zero rotation at asymptotic infinity. In particular, there is dragging, usually known as the Lense-Thirring effect, at infinity, i.e., $Z(r=\infty)\neq 0$. Thus, locally zero angular momentum observers (ZAMOs) [[39]39,,  [40]40] located at asymptotic infinity rotate with a non-zero constant angular velocity $\Omega=Z(r=\infty)$.
Although non-zero $\Omega$ can certainly lower (and raise) the mass, see e.g. the general expression (C.27) of the mass in the Appendix, it does not always result in negative mass solutions. A simple example is the BTZ solution with non-zero rotation at asymptotic infinity. To allow rotation at infinity we shift $Z$ in (4.29) by a non-zero constant. We denote the constant by $\overline{\lambda}$ and make the change
@@ 5. Discussion
The black string solutions are described by (3.15) which we again write below for convenience. They are obtained from ${\cal A}_{3}$ by applying the simple coordinate transformation (3.13) (and making the change of variable (3.14)). They carry the axion charge per unit length $Q$ (3.19). The black string solutions (labelled by the quartet $(\rho_{-},\rho_{+};\gamma;\lambda)$) are
@@
$\eqalign{e^{-2\phi}&={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}.}$ (5.4)5.4( 5.4 )
$\lambda$ is an arbitrary constant which determines the value of the $B$ field, i.e. $B_{tx}$, at $\rho=\infty$. It does not enter in $H$. However, it has a non-trivial implication.28 The black string solution is invariant under the combined transformations $\rho_{\pm}\to\rho_{\mp}$, $\gamma\to-\gamma$ and $\lambda\to-\lambda$.
The black string solution (5.3) has inner and outer horizons. This is also a general feature of rotating black holes. The inner horizon is at $\rho=\rho_{-}$ and the outer horizon is at $\rho=\rho_{+}$. In this paper, we have in general assumed, unless stated otherwise, $x$ is compact and $\rho$ is positive. In the case there is no curvature singularity we identify $x$ so that there is no conical singularity. Its period does not depend on the gauge parameter $\lambda$. To the best of my knowledge the black string solutions (5.3) have not been reported elsewhere in the literature.
For positive $\gamma$, i.e., $0<\gamma\leq 1$, the solution has a ring curvature singularity at $\rho=\rho_{0}\geq\rho_{+}$ (4.1). Therefore, the ring singularity cannot be continuously contracted or deformed to the origin without crossing the horizon(s). For $\gamma=1$ the outer horizon becomes singular, see (3.40). The singularity is behind the horizon in the interior region, i.e., $\rho_{0}>\rho_{-}$. Note the interior region is defined in the usual way as the part of the spacetime that contains the singularity. See the Penrose diagrams in Fig. 4 and Fig. 5.
@@
The Kalb-Ramond field is invariant under the Lorentz boost. The resulting metric however now has no cross term and thus is static, see (3.44). We subsequently applied T-duality transformation along the $\widehat{x}$ direction. This resulted in the black hole solution (3.51),
@@
To the best of my knowledge this black hole solution has not been reported elsewhere in the literature. The black hole solution has a singularity at $\rho=\rho_{-}$ and an event horizon at $\rho=\rho_{+}$. See the Penrose diagram in Fig. 3. Its (analogous) ADM mass and angular momentum are given (with $g_{s}=1$ and in units in which $\kappa^{2}=\pi$ and $\alpha^{\prime}=1$) by (4.38), (4.41)
@@
At $\gamma=-1$ we note that $J=0$ and the mass $M$ is independent of the gauge parameter $\lambda$. $M$ is determined only by the radii $\rho_{+}$ and $\rho_{-}$. At $\lambda=-1/2\gamma$, the angular velocity at infinity is zero, i.e., $\Omega=0$. In this case the ADM mass and angular momentum reduce to (4.43)
@@
In this paper we also noted that a solution with a non-zero asymptotic Kalb-Ramond field is equivalent to a solution with no Kalb-Ramond field and non-zero rotation at asymptotic infinity. In general, non-zero rotation allows negative mass solutions.
In an upcoming work we study the thermodynamics and partition functions of the black hole and black string solutions. We also study the causal structure of the solutions in detail.
In this paper we mainly presented black hole and black string solutions. In a future work we hope to study in greater detail the marginal operator that generates the deformation on the world-sheet.
@@
The scalar curvature of the solution (which is described by (5.13) and (5.14)) is $R=-4\xi^{2}\cdot\left(\rho_{+}^{2}/l^{2}\right)\cdot\left(1/|\rho^{2}-\rho^{2}_{-}|\right)$, where $\xi^{2}=1-\left(\rho_{-}/\rho_{+}\right)^{2}$. The Kretschmann scalar is $K=R^{2}$.29 In general, the equality $K=R^{2}$ is true only in special cases. Thus, the solution has the peculiar feature that it contains a ring curvature singularity at $\rho={\overline{\rho}}_{+}=\rho_{-}$ in between the inner and outer horizons. The inner horizon is at $\rho={\overline{\rho}}_{-}$. The outer horizon is at $\rho=\rho_{+}$. Note they are not related by sign reflection, i.e. ${\overline{\rho}}_{-}\neq-\rho_{+}$. The Kalb-Ramond field is a constant (field). It can be fixed to zero. The Penrose diagram of the solution is given in Fig. 6. It is obtained by combining Fig. 3 and Fig. 5.
Fig. 6: The plot depicts the global structure of the string solution described by (5.13) and (5.14). The solution has a singularity at $\rho=\rho_{-}={\overline{\rho}}_{+}$. It is represented by the zigzag lines. $\rho=0$ is a causal singularity. That is, the maximal analytic extension behind the surface $\rho=0$ contains closed time-like curves.
Note that usually the diagram ends at either side of $\rho={\overline{\rho}}_{+}=\rho_{-}$, however here, since there are no closed time-like curves, the diagram extends outward to infinity and inward to zero. Therefore, the maximal extension simply extends infinitely in both vertical directions.
We expect the black hole solution (described by (5.13) and (5.14)) to have a generalization to $(3+1)$ dimensions with similar feature. That is, with a curvature singularity in between the inner and outer horizons. We hope to study the corrections in $g_{s}$ and their roles in resolving the singularity in a future work. In relation to this, we also would like to better understand the solution (3.54),
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{0}\over r}\right)^{2}r^{2}}-\left(1-{r_{0}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{0}\over r}\right)}\left(d\chi-{r_{0}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{0}|)+{\tenrm const.}.}$ (5.15)5.15( 5.15 )
Its Ricci scalar is zero.